OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). If necessary, use the slash mark (/) for the division symbol. F(x)=(2x-3)/5 f^-1(y)=_______

OpenStudy (anonymous):

help please

hero (hero):

I showed you the steps already.

OpenStudy (anonymous):

i know but this is kind of hard for me

hero (hero):

The first step is y = F(x), right?

OpenStudy (anonymous):

yes Sr.

hero (hero):

Please, just call me Hero

hero (hero):

If y = F(x), then \[y = \frac{2x - 3}{5}\]

hero (hero):

The next step is to swap x and y

OpenStudy (anonymous):

ok

hero (hero):

Because you're basically taking the graph of the function and flipping it over the axis over the line y = x \[x = \frac{2y - 3}{5}\]

hero (hero):

That's why you swap x and y. But to complete the process of finding the inverse, you have to isolate y.

hero (hero):

\[5x = 2y - 3\] \[5x + 3 = 2y\] \[\frac{5x + 3}{2} = y\] This gives us the inverse of F(x)

hero (hero):

So we replace y with \(F^{-1}(x)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it looks somethin like 5x+3/2 right

hero (hero):

(5x + 3)/2

OpenStudy (anonymous):

do i have to pur () < this

hero (hero):

Yes, you do

hero (hero):

Otherwise it gets misinterpreted

OpenStudy (anonymous):

OMG so i basically got everythin wrongggggg >:|

hero (hero):

Well that depends

hero (hero):

Did you input it like this: \[\frac{5x + 3}{2}\] or like this: 5x + 3/2

OpenStudy (anonymous):

the sec one

hero (hero):

It should be (5x + 3)/2 if you write it horizontally